OpenStudy (anonymous):

Prove that sine squared theta - cosine squared theta = 2sin squared theta-1

Nnesha (nnesha):

\[\huge\rm sin^ 2 \theta - \cos^2 \theta = 2 \sin^2 \theta -1\] sin^2 = ?? cos^2 = ??